# Nontriviality of rings of integral-valued polynomials

### v1

@@ Abstract
MSC Primary 12J20, 13F30, 13A18, 13B25, 13F20.
@@ 1 Introduction and Motivation
Let $\overline{\mathbb{Q}}$ be a fixed algebraic closure of the field of rational numbers, and let $\overline{\mathbb{Z}}$ be the integral closure of $\mathbb{Z}$ in $\overline{\mathbb{Q}}$, which is the ring of all algebraic integers. For a subset $S\subseteq\overline{\mathbb{Z}}$, a polynomial $f\in\mathbb{Q}[X]$ is said to be integral-valued if $f(s)\in\overline{\mathbb{Z}}$ for all $s\in S$. Sets of integral-valued polynomials often form rings with interesting properties, and these rings have been of interest in recent years [7, 13, 15, 18, 20, 26].
@@
Note that $\textnormal{Int}_{\mathbb{Q}}(\mathcal{A}_{n})=\textnormal{Int}_{\mathbb{Q}}(\mathcal{A}_{n},\overline{\mathbb{Z}})$. It is stated in [15, p. 2482] that $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ lies properly between $\mathbb{Z}[X]$ and $\textnormal{Int}(\mathbb{Z})$ for every subset $S$ of $\overline{\mathbb{Z}}$ properly containing $\mathbb{Z}$. The containment $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})\subsetneq\textnormal{Int}(\mathbb{Z})$ holds whenever $\mathbb{Z}\subsetneq S$ (see [13]), but we demonstrate below in Example 1.3 that $\mathbb{Z}[X]$ may equal $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$. Hence, the aforementioned statement from [15, p. 2482] is false. [^3] Thankfully, the fallacious claim from [15] was merely an expository remark, and does not affect the results of that paper.
@@ 2 Nontriviality over Valuation Domains
Throughout, $V$ is a valuation domain with valuation $v$, value group $\Gamma_{v}$, maximal ideal $M$, and fraction field $K$. We begin by recalling the definitions of pseudo-monotone sequences and related objects, which were introduced in [6] in order to study the polynomial closure of subsets of a rank one valuation domain and studied further in [21, 23]. After proving some basic properties of such sequences, we use them to characterize when $\textnormal{Int}_{K}(S,V)$ is nontrivial (see Theorem 2.6).
@@
Let $\Lambda$ be an index set, and let $E=\{s_{i}\}_{i\in\Lambda}$ be a sequence in $K$. Then, $E$ is said to be
@@
  3. (iii)
    pseudo-stationary if $v(s_{i}-s_{j})=v(s_{k}-s_{\ell})$ for all $i\neq j\in\Lambda$, $k\neq\ell\in\Lambda$.
A sequence that satisfies any one of these three properties is said to be pseudo-monotone.
@@
    if $E$ is pseudo-convergent, for each $i\in\Lambda$ we set $\delta_{i}=v(s_{i}-s_{j})$, with $j\in\Lambda$ and $j>i$;
@@
  3. (iii)
    if $E$ is pseudo-stationary, we let $\delta=v(s_{i}-s_{j})$ for any $i,j\in\Lambda$, $i\neq j$.
@@
Each pseudo-monotone sequence $E=\{s_{i}\}_{i\in\Lambda}$ in $K$ has an associated breadth ideal, which we denote by $\textnormal{Br}(E)$. The definition of $\textnormal{Br}(E)$ depends on which one the three properties from Definition 2.1 that $E$ satisfies.
@@
    If $E$ is pseudo-convergent, then $\textnormal{Br}(E)=\{x\in K\mid v(x)>\delta_{i}\text{ for all }i\in\Lambda\}$;
@@
Note that if $E$ is pseudo-stationary and $\delta=v(c)$ for some $c\in K$, then $\textnormal{Br}(E)=cV$ is a principal (fractional) ideal. Also, if $E\subseteq V$ is any pseudo-monotone sequence, then $\textnormal{Br}(E)\subseteq V$, and $\textnormal{Br}(E)\subseteq M$ if $E$ is pseudo-divergent.
Finally, we recall that, given $a\in K$ and $\gamma\in\Gamma_{v}$, the ball of center $a$ and radius $\gamma$ is $B(a,\delta)=\{x\in K\mid v(x-a)\geq\delta\}$. For each $b\in B(a,\delta)$ it is well known that $B(a,\delta)=B(b,\delta)$.
@@
Note that $\delta=\infty$ in (1) corresponds to $b=0$ in (2). In this case, the set $S$ is finite.
@@
(2) $\Rightarrow$ (1) Assume that (2) holds. Let $t_{1},\ldots,t_{m}\in S$ be such that $S/bV=\{t_{i}+bV\mid 1\leq i\leq m\}$. Take $T=\{t_{1},\ldots,t_{m}\}$ and $\delta=v(b)>0$. Then, for any $s\in S$, there exists $i$ such that $s+bV=t_{i}+bV$, and for this $i$ we have $v(s-t_{i})\geq v(b)=\delta$. ∎
@@
The theorem is trivial if $S$ is finite (which is equivalent to having $\delta=\infty$ in (2), or $b=0$ in (3)), so we will assume throughout that $S$ is infinite.
@@
Next, let $L/K$ be the splitting field of $g(X)$, and let $w$ be an extension of $v$ to $L$. In $L$, we have $g(X)=a(X-\alpha_{1})\cdots(X-\alpha_{m})$ for some $\alpha_{1},\ldots,\alpha_{m}\in L$. Let $\delta=(v(d)-v(a))/m$, which is greater than 0. Then, for all $s\in S$, there exists $1\leq i\leq m$ such that $w(s-\alpha_{i})\geq\delta$. Indeed, if this is not the case, then for some $s\in S$,
$v(f(s))=w(f(s))<v(a)+m\delta-v(d)=0,$
@@
For each $1\leq i\leq m$, let $B_{w}(\alpha_{i},\delta)=\{x\in L\mid w(x-\alpha_{i})\geq\delta\}$ and let $B_{i}=B_{w}(\alpha_{i},\delta)\cap V$. Whenever $B_{i}$ has nonempty intersection with $S$, we may choose $t_{i}\in B_{i}\cap S$, and then $B_{i}=B(t_{i},\delta)$. Take $T=\{t_{i}\mid 1\leq i\leq m\text{ and }B_{i}\cap S\neq\varnothing\}$. Then (2) holds.
(2) $\Leftrightarrow$ (3) This is the content of Lemma 2.5.
@@
Choose $s_{1}\in S$. For each $n\geq 1$, choose—if possible—$s_{n+1}\in S$ such that $v(\prod_{i=1}^{n}(s_{n+1}-s_{i}))=0$. We cannot do this indefinitely, since then $\{s_{n}\}_{n\in\mathbb{N}}$ would be pseudo-stationary with breadth ideal equal to $V$. So, there exists $N\geq 1$ such that
$v(s_{i}-s_{j})=0\text{ for all }1\leq i<j\leq N,\text{ and }v(\textstyle\prod_{i=1}^{N}(s-s_{i}))>0\text{ for all }s\in S.$ (2.8)
@@
If we continue in this way, we can produce a sequence $E=\{s_{N+k}\}_{k\in\mathbb{N}}\subseteq S$ and a sequence $\{\delta_{k}\}_{k\in\mathbb{N}}\subseteq\{\gamma\in\Gamma_{v}\mid\gamma>0\}$ such that $\{\delta_{k}\}_{k\in\mathbb{N}}$ decreases to 0 and for each $k,\ell\in\mathbb{N}$, $k<\ell$ we have
@@ 3 Local Case
In this section, we show how Theorem 2.6 can be used to decide whether or not $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ is trivial. Let $\mathbb{P}$ be the set of all prime numbers. For $p\in\mathbb{P}$, $v_{p}$ is the standard $p$-adic valuation, $\mathbb{Z}_{p}$ denotes the ring of $p$-adic integers, $\mathbb{Q}_{p}$ is the field of $p$-adic numbers, and $\mathbf{F}_{p}$ is the finite field with $p$ elements. We let $\overline{\mathbb{Q}_{p}}$ be an algebraic closure of $\mathbb{Q}_{p}$, and let $\overline{\mathbb{Z}_{p}}$ be the absolute integral closure of $\overline{\mathbb{Z}_{p}}$. Note that $\overline{\mathbb{Z}_{p}}$ is a rank one non-discrete valuation domain with residue field equal to an algebraic closure of $\mathbb{F}_{p}$.
@@
Since $\overline{\mathbb{Z}_{p}}$ is a valuation domain with fraction field $\overline{\mathbb{Q}_{p}}$, combining Theorem 2.6 and Theorem 3.2 yields the following corollary.
@@
The proof of Theorem 3.2 is quite long and requires numerous intermediate results. To keep things organized, we will prove the equivalences in the theorem one at a time.
@@
First, we show that our rings of integer-valued polynomials are well-behaved with respect to localization at primes of $\mathbb{Z}$.
@@
Since ¯$\overline{\mathbb{Z}}=\bigcap_{p\in\mathbb{P}}\overline{\mathbb{Z}_{(p)}}$, we clearly have
@@
First, let let $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ and $n\in\mathbb{Z}\setminus p\mathbb{Z}$. Then, for each $s\in S$, we have that $\frac{f(s)}{n}$ is integral over $\mathbb{Z}_{(p)}$ by Lemma 3.4. Thus, $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})_{(p)}\subseteq\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})$.
The proof of the other containment is similar to the proof of [8, (7)]. Let $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})$. There exists $d\in\mathbb{Z},d\not=0$ such that $df(X)\in\mathbb{Z}[X]$. Suppose that $d=p^{a}t$, for some $a\geq 0$ and $t\in\mathbb{Z}$ not divisible by $p$. Let $g(X)=tf(X)$. Then $g\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})$ because $t$ is invertible in $\mathbb{Z}_{(p)}$. Moreover, for each prime $q\not=p$, $g\in\mathbb{Z}_{(q)}[X]\subseteq\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(q)}})$. Now, we have
@@
so that $f=\frac{g}{t}$ is in $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})_{(p)}$. ∎
@@
First, assume $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})$ and let $\alpha\in\Sigma_{p}(S)$. Then, there exists $\beta\in S$ such that $\alpha$ and $\beta$ have the same minimal polynomial over $\mathbb{Q}$. Moreover, $\alpha=\sigma(\beta^{\prime})$, where $\beta^{\prime}$ is a conjugate of $\beta$ over $\mathbb{Q}$ and $\sigma:\mathbb{Q}(\beta^{\prime})\to\mathbb{Q}_{p}(\alpha)$ is a $\mathbb{Q}$-embedding (which corresponds to some prime ideal $P$ of $O_{\mathbb{Q}(\beta^{\prime})}$ above $p$). By assumption, $f(\beta)$ is integral over $\mathbb{Z}_{(p)}$; but, $f$ is also integral-valued on $\beta^{\prime}$, even if $\beta^{\prime}$ does not lie in $S$. Hence, $f(\alpha)=\sigma(f(\beta^{\prime}))\in O_{\mathbb{Q}_{p}(\alpha)}\subseteq\overline{\mathbb{Z}_{p}}$.
@@
(1) $\Leftrightarrow$ (2) By [22, Lemma 2.20], the integral closure of $\textnormal{Int}_{\mathbb{Q}_{p}}(S,\overline{\mathbb{Z}_{p}})=\textnormal{Int}_{\mathbb{Q}_{p}}(G_{p}(S),\overline{\mathbb{Z}_{p}})$ in $\overline{\mathbb{Q}_{p}}(X)$ is equal to $\textnormal{Int}_{\overline{\mathbb{Q}_{p}}}(G_{p}(S),\overline{\mathbb{Z}_{p}})$. In particular, $\textnormal{Int}_{\mathbb{Q}_{p}}(S,\overline{\mathbb{Z}_{p}})$ strictly contains $\mathbb{Z}_{p}[X]$ if and only if $\textnormal{Int}_{\overline{\mathbb{Q}_{p}}}(G_{p}(S),\overline{\mathbb{Z}_{p}})$ strictly contains $\overline{\mathbb{Z}_{p}}[X]$.
@@
(3) $\Rightarrow$ (2) Assume $\textnormal{Int}_{\overline{\mathbb{Q}_{p}}}(S,\overline{\mathbb{Z}_{p}})$ is nontrivial. By Theorem 2.6, there exists a finite set $T\subseteq S$ and $\delta>0$ such that for each $s\in S$, there exists $t\in T$ with $v(s-t)\geq\delta$. Note that since $T$ is finite, so is the set $G_{p}(T)$ of all images of $T$ under $G_{p}$. Let $\sigma(s)\in G_{p}(S)$, where $\sigma\in G_{p}$ and $s\in S$. Then, there exists $t\in T$ with $v(s-t)\geq\delta$. Since ¯$\overline{\mathbb{Q}}_{p}$ is Henselian, we have $v(\sigma(\alpha))=v(\alpha)$ for all $\alpha\in\overline{\mathbb{Q}}_{p}$. So,

### v2

@@ Abstract
MSC Primary 12J20, 13F30, 13A18, 13B22, 13B25, 13F20.
Data Availability Statement
Data sharing is not applicable to this article. No new data was created or analyzed for this article.
Conflict of Interest
The authors declare that they have no conflicts of interest.
@@ 1 Introduction and Motivation
footnotetext: The second author has been partially supported by the University of Padova.
Let $\overline{\mathbb{Q}}$ be a fixed algebraic closure of the field of rational numbers, and let $\overline{\mathbb{Z}}$ be the integral closure of $\mathbb{Z}$ in $\overline{\mathbb{Q}}$, which is the ring of all algebraic integers. For a subset $S\subseteq\overline{\mathbb{Z}}$, a polynomial $f\in\mathbb{Q}[X]$ is said to be integral-valued if $f(s)\in\overline{\mathbb{Z}}$ for all $s\in S$. When $S\subseteq\mathbb{Z}$, the definition reduces to the classical notion of integer-valued polynomials, i.e., $f(S)\subseteq\mathbb{Z}$. We adopt the general term ‘integral-valued’ for polynomials with rational coefficients to underline the fact that the evaluation occurs at elements which are integral over $\mathbb{Z}$. Sets of integral-valued polynomials often form rings with interesting properties, and these rings have been of interest in recent years [7, 13, 15, 18, 21, 27].
@@
Note that $\textnormal{Int}_{\mathbb{Q}}(\mathcal{A}_{n})=\textnormal{Int}_{\mathbb{Q}}(\mathcal{A}_{n},\overline{\mathbb{Z}})$. It is stated in [15, p. 2482] that $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ lies properly between $\mathbb{Z}[X]$ and $\textnormal{Int}(\mathbb{Z})$ for every subset $S$ of $\overline{\mathbb{Z}}$ properly containing $\mathbb{Z}$. The containment $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})\subsetneq\textnormal{Int}(\mathbb{Z})$ holds whenever $\mathbb{Z}\subsetneq S$ (see [13]), but we demonstrate below in Example 1.3 that $\mathbb{Z}[X]$ may equal $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$. Hence, the aforementioned statement from [15, p. 2482] is false.Thankfully, the fallacious claim from [15] was merely an expository remark, and does not affect the results of that paper.
@@ 2 Nontriviality over Valuation Domains
Throughout, $V$ is a valuation domain with valuation $v$, value group $\Gamma_{v}$, maximal ideal $M$, and fraction field $K$. We begin by recalling the definitions of two kinds of pseudo-monotone sequences and related objects, which were introduced in [6] in order to study the polynomial closure of subsets of a rank one valuation domain and studied further in [22, 24]. These definitions originate from the notion of pseudo-convergent sequence given by Ostrowski in [20], and we will not use this kind of sequence in this paper. After proving some basic properties of such sequences, we use them to characterize when $\textnormal{Int}_{K}(S,V)$ is nontrivial (see Theorem 2.7).
@@
Let $\Lambda$ be an index set which we assume to be infinite and well-ordered, and let $E=\{s_{i}\}_{i\in\Lambda}$ be a sequence in $K$. Then, $E$ is said to be
@@
    pseudo-stationary if $v(s_{i}-s_{j})=v(s_{k}-s_{\ell})\in\Gamma_{v}$ for all $i\neq j\in\Lambda$, $k\neq\ell\in\Lambda$.
A sequence that satisfies either of these two properties is said to be pseudo-monotone.
@@
    if $E$ is pseudo-stationary, we let $\delta=v(s_{i}-s_{j})$ for any $\in\Lambda, iji\neq j$.
@@
Each pseudo-monotone sequence $E=\{s_{i}\}_{i\in\Lambda}$ in $K$ has an associated breadth ideal, which we denote by $\textnormal{Br}(E)$. The definition of $\textnormal{Br}(E)$ depends on which one of the two properties from Definition 2.1 that $E$ satisfies.
@@
Note that if $E$ is pseudo-stationary and $\delta=v(c)$ for some $c\in K$, then $\textnormal{Br}(E)=cV$ is a principal (fractional) ideal. Also, if $E\subset V$ is any pseudo-monotone sequence, then $\textnormal{Br}(E)\subseteq V$, and $\textnormal{Br}(E)\subseteq M$ if $E$ is pseudo-divergent.
Finally, we recall that, given $a\in K$ and $\delta\in\Gamma_{v}$, the ball of center $a$ and radius $\delta$ is $B(a,\delta)=\{x\in K\mid v(x-a)\geq\delta\}$. For each $b\in B(a,\delta)$ it is well known that $B(a,\delta)=B(b,\delta)$.
@@
For a value group $\Gamma_{v}$ associated to the valuation $v$, the divisible hull of $\Gamma_{v}$ is $\mathbb{Q}\Gamma_{v}:=\Gamma_{v}\otimes_{\mathbb{Z}}\mathbb{Q}$. Note that if $L/K$ is an algebraic extension and $w$ is an extension of $v$ to $L$, then $w(\alpha)\in\mathbb{Q}\Gamma_{v}$ for all $\alpha\in L$.
@@
  1. (1)
    There exist a finite subset $T\subseteq S$ and $\lambda\in\mathbb{Q}\Gamma_{v}\cup\{\infty\}$ such that $\lambda>0$ and for all $s\in S$, there exists $t\in T$ such that $v(s-t)\geq\lambda$.
@@
Note that $\lambda=\infty$ or $\delta=\infty$ in (1) or (2) corresponds to $b=0$ in (3). In this case, the set $S$ is finite. So, we assume throughout that $\lambda\neq\infty$ and $\delta\neq\infty$.
(1) $\Rightarrow$ (2) Assume that (1) holds. Let $\Gamma^{\prime}=\{\gamma\in\Gamma_{v}\mid\gamma\geq\lambda\}$. Suppose first that $\Gamma^{\prime}$ contains a minimum element $\gamma^{\prime}$. Then, for all $x\in V$, $v(x)\geq\lambda$ if and only if $v(x)\geq\gamma^{\prime}$, so we may take $\delta=\gamma^{\prime}$ and (2) holds. So, assume that $\Gamma^{\prime}$ has no minimum element. In particular, this means that $\lambda\notin\Gamma_{v}$.
Since $\lambda\in\mathbb{Q}\Gamma_{v}\setminus\Gamma_{v}$, there exists $\gamma\in\Gamma_{v}$ and $m\in\mathbb{Z}$, $m\geq 2$, such that $\lambda=\gamma/m$. For each $1\leq k\leq m-1$, let
$I_{k}=\{\varepsilon\in\Gamma^{\prime}\mid k\lambda<\varepsilon\leq(k+1)\lambda\}.$
We claim that there exists $k\in\{1,\ldots,m-1\}$ and $\varepsilon,\varepsilon^{\prime}\in\Gamma^{\prime}$ such that
$k\lambda<\varepsilon<\varepsilon^{\prime}\leq(k+1)\lambda.$ (2.6)
To see this, let $\delta_{1}=\gamma=m\lambda\in\Gamma^{\prime}$. Since $\Gamma^{\prime}$ has no minimum value in $\Gamma_{v}$, there exist $\delta_{2},\ldots,\delta_{m}\in\Gamma^{\prime}$ such that $\delta_{1}>\delta_{2}>\cdots>\delta_{m}>\lambda$. Since there are m1m-1 sets $I_{1},\ldots,I_{m-1}$, some $I_{k}$ must contain at least two distinct elements $\delta_{i},\delta_{j}\in\{\delta_{1},\ldots,\delta_{m}\}$ such that $\delta_{i}>\delta_{j}$. For this $k$, taking $\varepsilon=\delta_{j}$ and $\varepsilon^{\prime}=\delta_{i}$ satisfies (2.6). Moreover, (2.6) implies that $0<\varepsilon^{\prime}-\varepsilon<\lambda$. Thus, we may take $\delta=\varepsilon^{\prime}-\varepsilon\in\Gamma_{v}$, and then for $x\in V$, $v(x)\geq\lambda$ implies that $v(x)\geq\delta>0$. Hence, (2) holds.
@@
(3) $\Rightarrow$ (1) Assume that (3) holds. Let $t_{1},\ldots,t_{m}\in S$ be such that $S/bV=\{t_{i}+bV\mid 1\leq i\leq m\}$. Take $T=\{t_{1},\ldots,t_{m}\}$ and $\lambda=v(b)>0$. Then, for any $s\in S$, there exists $i$ such that $s+bV=t_{i}+bV$, and for this $i$ we have $v(s-t_{i})\geq v(b)=\lambda$.∎
@@
  2. (2)
    There exist a finite subset $T\subseteq S$ and $\lambda\in\mathbb{Q}\Gamma_{v}\cup\{\infty\}$ such that $\lambda>0$ and for all $s\in S$, there exists $t\in T$ such that $v(s-t)\geq\lambda$.
@@
The theorem is trivial if $S$ is finite (which is equivalent to having $\lambda=\infty$ in (2), $\delta=\infty$ in (3), or $b=0$ in (4)), so we will assume throughout that $S$ is infinite.
@@
Next, let $L/K$ be the splitting field of $g(X)$, and let $w$ be an extension of $v$ to $L$. In $L$, we have $g(X)=a(X-\alpha_{1})\cdots(X-\alpha_{m})$ for some $\alpha_{1},\ldots,\alpha_{m}\in L$. Let $\lambda=(v(d)-v(a))/m\in\mathbb{Q}\Gamma_{v}$, which is greater than 0. Then, for all $s\in S$, there exists $1\leq i\leq m$ such that $w(s-\alpha_{i})\geq\lambda$. Indeed, if this is not the case, then for some $s\in S$,
$v(f(s))=w(f(s))<v(a)+m\lambda-v(d)=0,$
@@
For each $1\leq i\leq m$, let $B_{w}(\alpha_{i},\lambda)=\{x\in L\mid w(x-\alpha_{i})\geq\lambda\}$ and let $B_{i}=B_{w}(\alpha_{i},\lambda)\cap S$. Whenever $B_{i}$ has nonempty intersection with $S$, we may choose $t_{i}\in B_{i}\cap S$. Take $T=\{t_{i}\mid 1\leq i\leq m\text{ and }B_{i}\cap S\neq\varnothing\}$. Note that $T\not$$\varnothing$ because we showed above that $S\subseteq\bigcup_{i=1}^{m}B_{w}(\alpha_{i},\lambda)$. Given $s\in S$, find $i$ such that $w(s-\alpha_{i})\geq\lambda$. Then, since $w(t_{i}-\alpha_{i})\geq\lambda$, we have $v(s-t_{i})=w(s-t_{i})\geq\lambda$. Thus, (2) holds.
(2) $\Leftrightarrow$ (3) $\Leftrightarrow$ (4) This is the content of Lemma 2.5.
@@
[…]
@@
If we continue in this way, we can produce a sequence $E=\{s_{N+k}\}_{k\in\mathbb{N}}\subseteq S$ and a sequence $\{\delta_{k}\}_{k\in\mathbb{N}}\subseteq\{\gamma\in\Gamma_{v}\mid\gamma>0\}$ such that $\{\delta_{k}\}_{k\in\mathbb{N}}$ decreases to 0 and for each $\ell\in\mathbb{N}, k<k$$\ell$ we have
@@ 3 Local Case
In this section, we show how Theorem 2.7 can be used to decide whether or not $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ is trivial. Let $\mathbb{P}$ be the set of all prime numbers. For $p\in\mathbb{P}$, $v_{p}$ is the standard $p$-adic valuation, $\mathbb{Z}_{(p)}$ is the localization of $\mathbb{Z}$ at $\mathbb{Z}, p\mathbb{Z}_{p}$ denotes the ring of $p$-adic integers, $\mathbb{Q}_{p}$ is the field of $p$-adic numbers, and $\mathbf{F}_{p}$ is the finite field with $p$ elements. We let $\overline{\mathbb{Q}_{p}}$ be an algebraic closure of $\mathbb{Q}_{p}$, and let $\overline{\mathbb{Z}_{p}}$ be the absolute integral closure of $\overline{\mathbb{Z}_{p}}$. Similarly, $\overline{\mathbb{Z}_{(p)}}$ denotes the integral closure of $\mathbb{Z}_{(p)}$ in $\overline{\mathbb{Q}}$. Note that $\overline{\mathbb{Z}_{p}}$ is a rank one non-discrete valuation domain with residue field equal to an algebraic closure of $\mathbb{F}_{p}$.
@@
The proof of Theorem 3.2 is quite long and requires numerous intermediate results. To keep things organized, we will prove the equivalences in the theorem one at a time. Before doing this, we mention the connection between Theorem 3.2 and our work from Section 2. Since $\overline{\mathbb{Z}_{p}}$ is a valuation domain with fraction field $\overline{\mathbb{Q}_{p}}$, combining Theorem 2.7 and Theorem 3.2 yields the following corollary.
@@
First, we show that our rings of integral-valued polynomials are well-behaved with respect to localization at primes of $\mathbb{Z}$.
@@
Since $\overline{\mathbb{Z}}=\bigcap_{p\in\mathbb{P}}\overline{\mathbb{Z}_{(p)}}$, we clearly have
@@
First, let $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ and $n\in\mathbb{Z}\setminus p\mathbb{Z}$. Then, for each $s\in S$, we have that $\frac{f(s)}{n}$ is integral over $\mathbb{Z}_{(p)}$ by Lemma 3.4. Thus, $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})_{(p)}\subseteq\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})$.
The proof of the other containment is similar to the proof of [8, (7)]. Let $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})$. There exists $d\in\mathbb{Z},d\not=0$ such that $df(X)\in\mathbb{Z}[X]$. Suppose that $d=p^{a}t$, for some $a\geq 0$ and $t\in\mathbb{Z}$ not divisible by $p$. Let $g(X)=tf(X)$. Then $g\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})$. Moreover, for each prime $q\not=p$, $g\in\mathbb{Z}_{(q)}[X]\subseteq\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(q)}})$. Now, we have
@@
so that $f=\frac{g}{t}$ is in $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})_{(p)}$ because $t$ is invertible in $\mathbb{Z}_{(p)}$. ∎
@@
First, assume $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})$ and let $\alpha\in\Sigma_{p}(S)$. Then, by definition of the set $\Sigma_{p}(S)$, there exists $\beta\in S$ such that $\alpha$ and $\beta$ have the same minimal polynomial over $\mathbb{Q}$. In particular, $\alpha=\sigma(\beta^{\prime})$, where $\beta^{\prime}\in\overline{\mathbb{Q}}$ is a conjugate of $\beta$ over $\mathbb{Q}$ and $\sigma:\mathbb{Q}(\beta^{\prime})\to\mathbb{Q}_{p}(\alpha)$ is a $\mathbb{Q}$-embedding (which corresponds to some prime ideal $P$ of $O_{\mathbb{Q}(\beta^{\prime})}$ above $p$). By assumption, $f(\beta)$ is integral over $\mathbb{Z}_{(p)}$; but, $f$ is also integral-valued on $\beta^{\prime}$, even if $\beta^{\prime}$ might not lie in $S$. Hence, $f(\alpha)=\sigma(f(\beta^{\prime}))\in O_{\mathbb{Q}_{p}(\alpha)}\subseteq\overline{\mathbb{Z}_{p}}$.
@@
(1) $\Leftrightarrow$ (2) By [23, Lemma 2.20], the integral closure of $\textnormal{Int}_{\mathbb{Q}_{p}}(S,\overline{\mathbb{Z}_{p}})=\textnormal{Int}_{\mathbb{Q}_{p}}(G_{p}(S),\overline{\mathbb{Z}_{p}})$ in $\overline{\mathbb{Q}_{p}}(X)$ is equal to $\textnormal{Int}_{\overline{\mathbb{Q}_{p}}}(G_{p}(S),\overline{\mathbb{Z}_{p}})$. In particular, $\textnormal{Int}_{\mathbb{Q}_{p}}(S,\overline{\mathbb{Z}_{p}})$ strictly contains $\mathbb{Z}_{p}[X]$ if and only if $\textnormal{Int}_{\overline{\mathbb{Q}_{p}}}(G_{p}(S),\overline{\mathbb{Z}_{p}})$ strictly contains $\overline{\mathbb{Z}_{p}}[X]$.
@@
(3) $\Rightarrow$ (2) Assume $\textnormal{Int}_{\overline{\mathbb{Q}_{p}}}(S,\overline{\mathbb{Z}_{p}})$ is nontrivial. By Theorem 2.7, there exists a finite set $T\subseteq S$ and $\delta>0$ such that for each $s\in S$, there exists $t\in T$ with $v(s-t)\geq\delta$. Note that since $T$ is finite, so is the set $G_{p}(T)$ of all images of $T$ under $G_{p}$. Let $\sigma(s)\in G_{p}(S)$, where $\sigma\in G_{p}$ and $s\in S$. Then, there exists $t\in T$ with $v(s-t)\geq\delta$. Since $\overline{\mathbb{Q}}_{p}$ is Henselian, we have $v(\sigma(\alpha))=v(\alpha)$ for all $\alpha\in\overline{\mathbb{Q}}_{p}$. So,